Select the x-coordinate of the vertex of the parabola defined by the function f(x) = -7x2 + 9x + 8.
Answer by stanbon(75887)

Select the x-coordinate of the vertex of the parabola defined by the function
f(x) = -7x2 + 9x + 8
----
Vertex occurs where x = -b/(2a) = -9/(2*-7) = 9/14
-----------------------------
Cheers,
Stan H.
=============================